Polarizer and analyzer are set with their polarizing directions parallel so that the intensity of transmitted light is maximum. Through what angle should either be turned so that the intensity is reduced to 80 % of the maximum intensity?

1
Expert's answer
2022-02-17T11:45:27-0500

Let the initial intensity =I

and final intensity "=\frac{4I}{5}"

We know that

"I=I_o\cos^2 \theta"

"\frac{4I}{5}=I\cos^2 \theta"

"\cos^2\theta=\frac{4}{5}"

"\cos\theta=\frac{2}{\sqrt{5}}"

"\theta=cos^{-1} =26.6"

"\theta=26.6°"